# Statistical study for binary star evolution in dense embedded clusters

### v1

@@ 1 Introduction
A key factor in dynamical modelling of stellar population is the characterization of the binary population (BP), which comprises the distributions of binary parameters such as primary mass, orbital period, eccentricity, and mass ratio. Numerous studies have shown that the dynamical evolution of star clusters is significantly influenced by the presence and properties of binary systems (e.g., Heggie, 1975; Kroupa, 1995a, b; Parker et al., 2009; Kroupa and Petr-Gotzens, 2011). In the context of stellar population synthesis, particularly when accounting for unresolved binaries, a reliable model for the BP is crucial for accurately reconstructing the stellar mass function (e.g., Kroupa et al., 1991; Marks and Kroupa, 2011; Pang et al., 2024; Kroupa, 2025).
It is widely accepted that the Galactic field stellar population originates predominantly from the dissolution of embedded star clusters that form in molecular cloud clumps (e.g., Kroupa, 1995a, b; Lada and Lada, 2003; Bressert et al., 2010; Lada, 2010; Kroupa, 2011; Marks and Kroupa, 2011; Kroupa, 2025). Following the expulsion of residual gas, an embedded cluster rapidly loses its dynamical equilibrium, leading to the loss of a significant fraction of single and multiple stellar systems into the Galactic field. As shown in studies such as Kroupa et al. (2001) and Boily and Kroupa (2002), the cluster undergoes a phase of rapid expansion followed by re-virialization, ultimately reaching a gas-free state. These post-gas-expulsion phases are less dynamically active than the deeply embedded phase, suggesting that the BP’s dynamical evolution prior to gas expulsion warrants detailed investigation.
Direct observation of the BP, however, is challenging. Prior to gas expulsion, the residual gas and dust obscures stellar light, making direct observation nearly impossible. During the rapid expansion phase, stars become observable, and several techniques may be employed. For close binaries with semi-major axes $a<10$ AU, spectroscopic monitoring and eclipsing binary analysis can be used to infer binary properties (e.g., Sana et al., 2012; Kounkel et al., 2019), though these methods require long observational times and are less sensitive to low mass ratios. Interferometry, sparse aperture masking, high-resolution imaging, and astrometry can detect binaries at intermediate separations ($a\approx 1$–300 AU), depending on the distance to the system (e.g., Sana et al., 2014; De Furio et al., 2019). However, these techniques are often infeasible for distant star-forming regions. For systems with high mass ratios or wide separations, analyses using colour–magnitude diagrams (CMDs) and proper motion data are effective (e.g., Tokovinin and Briceño, 2020; Jadhav et al., 2021), although such methods also have clear limitations. As a result, detailed modelling of the BP using $N$-body simulations becomes essential.
@@
Modelling the dynamical evolution of binary-rich, high-mass star-forming regions such as the Orion Nebula Cluster (ONC), which contains on the order of $10^{4}$ stars and brown dwarfs, remains a major challenge for direct $N$-body simulations. Moreover, the deeply embedded phase of open clusters likely had significantly higher masses than observed today. For instance, Kroupa et al. (2001) (see also Safaei et al., 2025) suggest that the initial mass in stars and brown dwarfs of the Pleiades was between 3000 and 4000 $\mathrm{M_{\odot}}$, compared to its present-day mass of roughly 800 $\mathrm{M_{\odot}}$. Since the dynamical evolution of binary systems is primarily driven by close encounters rather than by the overall cluster potential, a direct $N$-body method needs to be employed to quantify the dynamical processing of the BP (see, e.g., Spurzem and Kamlah, 2023).
@@ 2.1.1 Initial mass function
For the upper mass limit $m_{\rm max}$, we adopt the empirical $m_{\rm max}$$M_{\rm ecl}$ relation from Weidner and Kroupa (2006) and Pflamm-Altenburg and Kroupa (2007) (see also Yan et al., 2023) giving $m_{\rm max}=89.73\penalty 10000\ {\rm M_{\odot}}$. To suppress stochastic variations in the number of massive stars—which can significantly affect the dynamical evolution (see e.g. Wang and Jerabkova, 2021)—we use the optimal sampling method introduced by Kroupa et al. (2013). In conventional random sampling, a uniform random variable $X\in(0,1]$ is drawn, and the corresponding stellar mass $m_{\star}$ is found by solving
@@ 2.1.3 Initial phase-space distribution
We note that our setup assumes the embedded cluster to be initially in a state of virial equilibrium. Recent studies on BPs have explored non-equilibrium initial conditions, such as fractal structures (e.g. Dorval et al., 2017), showing that the environment can significantly influence the dynamical evolution of BPs. However, as pointed out by Kroupa et al. (2001), the initial conditions of embedded clusters are subject to physical constraints. In reality, star formation proceeds over about a Myr until gas expulsion. Once a star forms, it either falls toward the cluster’s potential minimum or is already on a ”virial-equilibrium orbit” (see also Moeckel and Clarke, 2011). In either case, the accumulating stellar system is continuously virializing as new stars are added, since the crossing timescale of an embedded cluster (see section 4.1) is much shorter than its formation timescale. Therefore, a young embedded cluster is most likely to be close to virial equilibrium.
@@ 2.2 The $N$-body integration
PeTar combines several advanced computational techniques to handle interactions across different distance scales: the Barnes-Hut tree algorithm (Barnes and Hut, 1986) for long-range interactions, a fourth-order Hermite integrator with block time-step (e.g., Aarseth, 2003) for intermediate-range interactions, and the slow-down algorithmic regularization method (Wang et al., 2020c) for short-range interactions.
@@
Gas is modelled as a static Plummer potential in the cluster-centric frame (see e.g., Kroupa et al., 2001; Banerjee and Kroupa, 2013). The gas mass is set to the mass of the embedded cluster divided by the star formation efficiency (SFE). Following the previous $N$-body modelling and to reduce the number of parameters, we assume here $\text{SFE}=1/3$. Both the stellar and gas potentials share the same Plummer radius. As discussed in Kroupa et al. (2001), the gas expulsion starts with a delay after the star formation phase, which is typically 0.6 Myr. Therefore we chose the end time of all simulations to be 0.6 Myr, because the thereafter emerging cluster expansion largely freezes the binary population. Gas expulsion drives substantial cluster expansion, the dynamical environment becomes less favourable for further binary processing beyond this point. Thus, limiting the integration time to 0.6 Myr allows us to isolate the effects of early cluster dynamics prior to the onset of gas-driven expansion.
We acknowledge that this model omits detailed star–gas interactions, such as dynamical friction with the gas and feedback-driven gas removal. In particular, gas dynamical friction may influence the evolution of high-mass binaries, potentially leading to tighter configurations or enhanced disruption rates. Recent studies have shown that gas interactions can harden binaries and extend the traditional soft–hard boundary (Rozner and Perets, 2024), while gas-assisted capture provides an efficient channel for the formation of new binaries, especially among massive stars in gas-rich environments (Rozner et al., 2023). Such effects are especially pronounced for massive or compact binaries. However, high-mass binaries constitute a minor fraction of the population in this study. As shown in Section 4.4, most dynamical modifications occur among low-mass binaries. Therefore, we conclude that our gas model is sufficient to capture the dominant dynamical effects relevant to the scope of this work.
@@ 4 Results and discussion
Following Casertano and Hut (1985), we define the local density for each star as the mass density within a sphere that includes its six nearest neighbours. Stars with local densities below $1\penalty 10000\ \mathrm{M_{\odot}}\penalty 10000\ \mathrm{pc}^{-3}$ are excluded in the calculation of the cluster’s centre and the boundary of the cluster . The cluster centre is then computed as the local-mass-density-weighted average position of the remaining stars, and the cut-off radius $R_{\rm cut}$ is defined as the distance from this centre to the furthest remaining star. This approach assumes that, under isotropy, the local density reflects the global density profile. In our simulations, $R_{\rm cut}$ typically lies between 2 and 3 pc. It is determined from the final snapshot (at 0.6 Myr), and the same value is used for earlier snapshots, with the cluster centre recalculated accordingly. During the evolution, stars are ejected from the cluster due to close encounters and the cluster expands slightly (see section 4.1 below). Thus $R_{\rm cut}$ from the final snapshot is large enough to cover all dynamical active systems during the calculation. In the analysis of binary populations, we only consider stars within $R_{\rm cut}$. At 0 Myr i.e. the initial time, there are 30 to 50 stars out of $R_{\rm cut}$ for each sample. However this will not affect the results, because we introduce no correlation between position and binary parameters in the initial setup. The half-mass radius and related quantities are determined using all stars, including those located beyond $R_{\rm cut}$, with respect to the updated cluster centre.
@@ 4.1 Timescale relevant to binary evolution
In our analysis, $t_{\rm cr}$ is calculated using the initial values of $R_{\rm h,m}$ and $\sigma_{\rm ecl}$. In our model, the gas potential increases the velocity dispersion, so we compute $t_{\rm cr}$ directly from Equation (6) with the velocity dispersion being calculated using the centre-of-mass velocities of systems within the half-mass radius. We also show here the half-mass two-body relaxation time $t_{\rm rh}$. The definition follows from e.g. Meylan and Heggie (1997),
@@
where $R_{\rm vir}$ is the virial radius of the cluster, in our case (Plummer sphere) $R_{\rm h,m}\approx 0.77R_{\rm vir}$, $N$ is the number of systems within $R_{\rm h,m}$ and $\operatorname{ln}\Lambda$ is the Coulomb logarithm. We use $\Lambda=0.02N$ based on the measurement of Giersz and Heggie (1996).
@@ 4.2 Dynamical evolution of binding energy and orbital period
Each binary is assigned a unique identifier (as implemented in the PeTar analysis tool). A binary present at both 0 and 0.6 Myr is classified as a surviving binary. Figure 4 shows the survival fraction at 0.6 Myr, defined as the number of surviving systems divided by the total number of binaries. The result indicates that more than 99.6% of binaries are survivors, implying that binary formation during this period is negligible, as also shown by Kroupa and Burkert (2001). More complex processes, such as component exchange (where one member of a binary is replaced during an encounter) or disruption followed by subsequent re-formation, are subsumed under the formation channel in this work. These events are difficult to track in the simulations because resolving them would require output time intervals much shorter than the typical orbital periods of binaries. However, since fewer than 0.7% of binaries at 0.6 Myr are non-surviving systems, their impact on the overall results is negligible.

### v2

@@ 1 Introduction
A key factor in dynamical modelling of stellar population is the characterization of the binary population (BP), which comprises the distributions of binary parameters such as primary mass, orbital period, eccentricity, and mass ratio. Numerous studies have shown that the dynamical evolution of star clusters is significantly influenced by the presence and properties of binary systems (e.g., Heggie, 1975; Kroupa, 1995a, b; Parker et al., 2009; Kroupa & Petr-Gotzens, 2011). In the context of stellar population synthesis, particularly when accounting for unresolved binaries, a reliable model for the BP is crucial for accurately reconstructing the stellar mass function (e.g., Kroupa et al., 1991; Marks & Kroupa, 2011; Pang et al., 2024; Kroupa, 2025).
It is widely accepted that the Galactic field stellar population originates predominantly from the dissolution of embedded star clusters that form in molecular cloud clumps (e.g., Kroupa, 1995a, b; Lada & Lada, 2003; Bressert et al., 2010; Lada, 2010; Kroupa, 2011; Marks & Kroupa, 2011; Kroupa, 2025). Following the expulsion of residual gas, an embedded cluster rapidly loses its dynamical equilibrium, leading to the loss of a significant fraction of single and multiple stellar systems into the Galactic field. As shown in studies such as Kroupa et al. (2001) and Boily & Kroupa (2002), the cluster undergoes a phase of rapid expansion followed by re-virialization, ultimately reaching a gas-free state. These post-gas-expulsion phases are less dynamically active than the deeply embedded phase, suggesting that the BP’s dynamical evolution prior to gas expulsion warrants detailed investigation.
Direct observation of the BP, however, is challenging. Prior to gas expulsion, the residual gas and dust obscures stellar light, making direct observation nearly impossible. During the rapid expansion phase, stars become observable, and several techniques may be employed. For close binaries with semi-major axes $a<10$ AU, spectroscopic monitoring and eclipsing binary analysis can be used to infer binary properties (e.g., Sana et al., 2012; Kounkel et al., 2019), though these methods require long observational times and are less sensitive to low mass ratios. Interferometry, sparse aperture masking, high-resolution imaging, and astrometry can detect binaries at intermediate separations ($a\approx 1$–300 AU), depending on the distance to the system (e.g., Sana et al., 2014; De Furio et al., 2019). However, these techniques are often infeasible for distant star-forming regions. For systems with high mass ratios or wide separations, analyses using colour–magnitude diagrams (CMDs) and proper motion data are effective (e.g., Tokovinin & Briceño, 2020; Jadhav et al., 2021), although such methods also have clear limitations. As a result, detailed modelling of the BP using $N$-body simulations becomes essential.
@@
Modelling the dynamical evolution of binary-rich, high-mass star-forming regions such as the Orion Nebula Cluster (ONC), which contains on the order of $10^{4}$ stars and brown dwarfs, remains a major challenge for direct $N$-body simulations. Moreover, the deeply embedded phase of open clusters likely had significantly higher masses than observed today. For instance, Kroupa et al. (2001) (see also Safaei et al., 2025) suggest that the initial mass in stars and brown dwarfs of the Pleiades was between 3000 and 4000 $\mathrm{M_{\odot}}$, compared to its present-day mass of roughly 800 $\mathrm{M_{\odot}}$. Since the dynamical evolution of binary systems is primarily driven by close encounters rather than by the overall cluster potential, a direct $N$-body method needs to be employed to quantify the dynamical processing of the BP (see, e.g., Spurzem & Kamlah, 2023).
@@ 2.1.1 Initial mass function
For the upper mass limit $m_{\rm max}$, we adopt the empirical $m_{\rm max}$$M_{\rm ecl}$ relation from Weidner & Kroupa (2006) and Pflamm-Altenburg & Kroupa (2007) (see also Yan et al., 2023) giving $m_{\rm max}=89.73\penalty 10000\ {\rm M_{\odot}}$. To suppress stochastic variations in the number of massive stars—which can significantly affect the dynamical evolution (see e.g. Wang & Jerabkova, 2021)—we use the optimal sampling method introduced by Kroupa et al. (2013). In conventional random sampling, a uniform random variable $X\in(0,1]$ is drawn, and the corresponding stellar mass $m_{\star}$ is found by solving
@@ 2.1.3 Initial phase-space distribution
We note that our setup assumes the embedded cluster to be initially in a state of virial equilibrium. Recent studies on BPs have explored non-equilibrium initial conditions, such as fractal structures (e.g. Dorval et al., 2017), showing that the environment can significantly influence the dynamical evolution of BPs. However, as pointed out by Kroupa et al. (2001), the initial conditions of embedded clusters are subject to physical constraints. In reality, star formation proceeds over about a Myr until gas expulsion. Once a star forms, it either falls toward the cluster’s potential minimum or is already on a ”virial-equilibrium orbit” (see also Moeckel & Clarke, 2011). In either case, the accumulating stellar system is continuously virializing as new stars are added, since the crossing timescale of an embedded cluster (see section 4.1) is much shorter than its formation timescale. Therefore, a young embedded cluster is most likely to be close to virial equilibrium.
@@ 2.2 The $N$-body integration
PeTar combines several advanced computational techniques to handle interactions across different distance scales: the Barnes-Hut tree algorithm (Barnes & Hut, 1986) for long-range interactions, a fourth-order Hermite integrator with block time-step (e.g., Aarseth, 2003) for intermediate-range interactions, and the slow-down algorithmic regularization method (Wang et al., 2020c) for short-range interactions.
@@
Gas is modelled as a static Plummer potential in the cluster-centric frame (see e.g., Kroupa et al., 2001; Banerjee & Kroupa, 2013). The gas mass is set to the mass of the embedded cluster divided by the star formation efficiency (SFE). Following the previous $N$-body modelling and to reduce the number of parameters, we assume here $\text{SFE}=1/3$. Both the stellar and gas potentials share the same Plummer radius. As discussed in Kroupa et al. (2001), the gas expulsion starts with a delay after the star formation phase, which is typically 0.6 Myr. Therefore we chose the end time of all simulations to be 0.6 Myr, because the thereafter emerging cluster expansion largely freezes the binary population. Gas expulsion drives substantial cluster expansion, the dynamical environment becomes less favourable for further binary processing beyond this point. Thus, limiting the integration time to 0.6 Myr allows us to isolate the effects of early cluster dynamics prior to the onset of gas-driven expansion.
We acknowledge that this model omits detailed star–gas interactions, such as dynamical friction with the gas and feedback-driven gas removal. In particular, gas dynamical friction may influence the evolution of high-mass binaries, potentially leading to tighter configurations or enhanced disruption rates. Recent studies have shown that gas interactions can harden binaries and extend the traditional soft–hard boundary (Rozner & Perets, 2024), while gas-assisted capture provides an efficient channel for the formation of new binaries, especially among massive stars in gas-rich environments (Rozner et al., 2023). Such effects are especially pronounced for massive or compact binaries. However, high-mass binaries constitute a minor fraction of the population in this study. As shown in Section 4.4, most dynamical modifications occur among low-mass binaries. Therefore, we conclude that our gas model is sufficient to capture the dominant dynamical effects relevant to the scope of this work.
@@ 4 Results and discussion
Following Casertano & Hut (1985), we define the local density for each star as the mass density within a sphere that includes its six nearest neighbours. Stars with local densities below $1\penalty 10000\ \mathrm{M_{\odot}}\penalty 10000\ \mathrm{pc}^{-3}$ are excluded in the calculation of the cluster’s centre and the boundary of the cluster . The cluster centre is then computed as the local-mass-density-weighted average position of the remaining stars, and the cut-off radius $R_{\rm cut}$ is defined as the distance from this centre to the furthest remaining star. This approach assumes that, under isotropy, the local density reflects the global density profile. In our simulations, $R_{\rm cut}$ typically lies between 2 and 3 pc. It is determined from the final snapshot (at 0.6 Myr), and the same value is used for earlier snapshots, with the cluster centre recalculated accordingly. During the evolution, stars are ejected from the cluster due to close encounters and the cluster expands slightly (see section 4.1 below). Thus $R_{\rm cut}$ from the final snapshot is large enough to cover all dynamical active systems during the calculation. In the analysis of binary populations, we only consider stars within $R_{\rm cut}$. At 0 Myr i.e. the initial time, there are 30 to 50 stars out of $R_{\rm cut}$ for each sample. However this will not affect the results, because we introduce no correlation between position and binary parameters in the initial setup. The half-mass radius and related quantities are determined using all stars, including those located beyond $R_{\rm cut}$, with respect to the updated cluster centre.
@@ 4.1 Timescale relevant to binary evolution
In our analysis, $t_{\rm cr}$ is calculated using the initial values of $R_{\rm h,m}$ and $\sigma_{\rm ecl}$. In our model, the gas potential increases the velocity dispersion, so we compute $t_{\rm cr}$ directly from Equation (6) with the velocity dispersion being calculated using the centre-of-mass velocities of systems within the half-mass radius. We also show here the half-mass two-body relaxation time $t_{\rm rh}$. The definition follows from e.g. Meylan & Heggie (1997),
@@
where $R_{\rm vir}$ is the virial radius of the cluster, in our case (Plummer sphere) $R_{\rm h,m}\approx 0.77R_{\rm vir}$, $N$ is the number of systems within $R_{\rm h,m}$ and $\operatorname{ln}\Lambda$ is the Coulomb logarithm. We use $\Lambda=0.02N$ based on the measurement of Giersz & Heggie (1996).
@@ 4.2 Dynamical evolution of binding energy and orbital period
Each binary is assigned a unique identifier (as implemented in the PeTar analysis tool). A binary present at both 0 and 0.6 Myr is classified as a surviving binary. Figure 4 shows the survival fraction at 0.6 Myr, defined as the number of surviving systems divided by the total number of binaries. The result indicates that more than 99.6% of binaries are survivors, implying that binary formation during this period is negligible, as also shown by Kroupa & Burkert (2001). More complex processes, such as component exchange (where one member of a binary is replaced during an encounter) or disruption followed by subsequent re-formation, are subsumed under the formation channel in this work. These events are difficult to track in the simulations because resolving them would require output time intervals much shorter than the typical orbital periods of binaries. However, since fewer than 0.7% of binaries at 0.6 Myr are non-surviving systems, their impact on the overall results is negligible.